OpenStudy (anonymous):

divide the polynomials. -3y^3 + 12y - 6 over -3

OpenStudy (lalaly):

\[\large{\frac{-3y^3+12y-6}{-3}} = \frac{\cancel{-3}y^3}{\cancel{-3}}+\frac{12y}{-3}-\frac{6}{-3}\]

OpenStudy (lalaly):

\[=\large{y^3-4y+2}\]

OpenStudy (anonymous):

thank you so much.

OpenStudy (lalaly):

=)